OpenStudy (anonymous):

pls help find the perimeter of rectangle if three of its vertices are (5,-2), (-3,-2), (-3,3)

OpenStudy (zale101):

perimeter of rectangle = 2 * sum of sides

OpenStudy (zale101):

x axis is (5 - (-3)) = 8. y axis (3 - (-2)) = 5. 8+5=13 so, 2 * sum of sides is 2 * 13 = 26

OpenStudy (anonymous):

thank you for your help.

OpenStudy (zale101):

 the length is 8 because 5 to -3 is extended so 5-(-3)=8 (the x-axis) the width is 5 because-3 to 2 is extended so 2-(-3)=5 (the x-axis) therefore, the perimeter is 8+5+8+5=26

OpenStudy (zale101):

2-(-3)=5 is in the y-axis**